If the angle between the two straight lines represented by 2x2+5xy+3y2+7y-98=0 is tan-1m then find m. 

Grade:11

1 Answers

Sunil Raikwar
9 years ago
required angle is tan-1{2squareroot(h2-ab)/a+b}=tan-1{2squareroot(25/4-6)/9}
=tan-1(1/9)
so m=1/9
Thanks & Regards
Sunil Raikwar
askIITians Faculty